OpenStudy (anonymous):

What is the solution to the equation? b + 0.8 = –1 b = @KailaMorgan7777 @dumbcow

OpenStudy (anonymous):

@KailaMorgan7777 @dumbcow @tom982

OpenStudy (anonymous):

b+0.8 = -1 b = -1-0.8 b = -1.8

OpenStudy (anonymous):

so it is -1.8 for the asnwer?

jimthompson5910 (jim_thompson5910):

tom982 is correct. Notice how he subtracted 0.8 from both sides. This is to undo the +0.8 done to the b on the left which isolates b.

OpenStudy (anonymous):

yes hes right!

OpenStudy (anonymous):

Thanks @jim_thompson5910, I should have explained that step or included another to show the additive inverse.

jimthompson5910 (jim_thompson5910):

That works too. Adding -0.8, the additive inverse of 0.8, to both sides also isolates b.